Grade 12Integral Calculus

evaluate
∫ 1 /((a)^x (b)^x) dx

taniska
12 Years agoGrade 12

1 Answer

Jitender Singh
12 Years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \frac{1}{a^{x}b^{x}}dx
I = \int \frac{1}{(ab)^{x}}dx
I = \int (ab)^{-x}dx
t = -x
dt = -dx
I = -\int (ab)^{t}dt
I = \frac{-(ab)^{t}}{log(ab)} + constant
I = \frac{-(ab)^{-x}}{log(ab)} + constant
I = \frac{-(ab)^{-x}}{log(a)+log(b)} + constant